### v1

# SoundMind: RL-Incentivized Logic Reasoning for Audio-Language Models

Abstract
While large language models have shown reasoning capabilities, their application to the audio modality, particularly in large audio-language models (ALMs), remains significantly underdeveloped. Addressing this gap requires a systematic approach, involving a capable base model, high-quality reasoning-oriented audio data, and effective training algorithms. In this study, we present a comprehensive solution: we introduce the Audio Logical Reasoning (ALR) dataset, consisting of 6,446 text-audio annotated samples specifically designed for complex reasoning tasks. Building on this resource, we propose SoundMind, a rule-based reinforcement learning (RL) algorithm tailored to endow ALMs with deep bimodal reasoning abilities. By training Qwen2.5-Omni-7B on the ALR dataset using SoundMind, our approach achieves state-of-the-art performance in audio logical reasoning. This work highlights the impact of combining high-quality, reasoning-focused datasets with specialized RL techniques, advancing the frontier of auditory intelligence in language models. Our code and the proposed dataset are available at https://github.com/xid32/SoundMind.
SoundMind: RL-Incentivized Logic Reasoning for Audio-Language Models
Xingjian Diao1, Chunhui Zhang1, Keyi Kong2, Weiyi Wu1, Chiyu Ma1, Zhongyu Ouyang1, Peijun Qing1, Soroush Vosoughi1, Jiang Gui1 1Dartmouth College, 2Shandong University xingjian.diao.gr@dartmouth.edu
1 Introduction
Figure 1: Overview of the end-to-end bimodal Natural Language Inference (NLI) task enabled by large audio-language models. The model receives input and produces output in either pure text or pure audio format.
Large language models (LLMs) have made remarkable strides in reasoning capabilities through innovations such as Chain-of-Thought (CoT) prompting and specialized reasoning architectures. Recent models such as OpenAI’s o1 model (Jaech et al., 2024) and Deepseek-R1 (Guo et al., 2025) have demonstrated exceptional performance on complex mathematical and programming tasks (Zhao et al., 2024; Team et al., 2025; Muennighoff et al., 2025). Particularly noteworthy is Deepseek-R1’s rule-based reinforcement learning approach, which enables emergent reasoning without relying on traditional frameworks like Monte Carlo Tree Search (Wan et al., 2024) or process reward models (Lightman et al., 2023). This general reasoning paradigm has successfully extended to the visual domain, where frameworks like Visual-CoT (Shao et al., 2024) have significantly enhanced multimodal models’ cognitive abilities for image and video reasoning.
Despite these advances in text and visual reasoning, a significant gap exists in audio reasoning and generation capabilities. While audio-language models (ALMs) like Audio Flamingo (Kong et al., 2024) and Qwen2.5-Omni-7B (Xu et al., 2025) have made progress in audio understanding, end-to-end audio reasoning remains underdeveloped. This limitation stems primarily from two factors: (1) the simplicity of existing audio reasoning datasets (Suzgun et al., 2023; Kong et al., 2024), which often contain only brief textual labels without proper audio modality annotations, and (2) the technical challenges in maintaining reasoning coherence during long-duration audio generation. Current CoT methods applied to audio often lead to hallucinations and performance degradation when generating extended reasoning sequences. The lack of aligned audio-text annotations further impedes research on reasoning-driven audio generation, creating a bottleneck in developing ALMs with sophisticated reasoning capabilities.
To address these challenges in audio reasoning (Figure 1), we introduce a comprehensive approach with two key components. First, we construct the Audio Logical Reasoning (ALR) dataset (Liu et al., 2023), containing 6,446 dual-modality samples with user content, CoT reasoning steps, final answers, and corresponding audio input/output. Derived from the Logi-QA 2.0-NLI dataset, ALR provides complete logical reasoning tasks with both content and response in text and audio modalities. Second, we propose the SoundMind algorithm, which builds upon the ALR dataset to address performance issues in long-form reasoning audio generation. Drawing inspiration from the Logic-RL framework (Xie et al., 2025a), SoundMind incorporates a strict format-based reward mechanism to prevent shortcut reasoning biased toward the textual modality. It leverages the REINFORCE++ algorithm (Hu, 2025) alongside the reward design principles from Deepseek-R1 for effective post-training. Our main contributions are as follows:
  • ALR Dataset: We release a high-quality dual-modality dataset containing 6,446 samples with user content, CoT reasoning, answers, and corresponding audio input/output. The ALR dataset includes deep reasoning annotations in both text and audio formats, providing a valuable resource for RL-based training of audio-language models.
  • SoundMind RL Algorithm: Taking into account the unique challenges of audio reasoning generation, we design a strict format-based reward mechanism that maintains reasoning coherence across modalities. By combining the improved REINFORCE++ algorithm with the reward design principles of Deepseek-R1, we fine-tune Qwen2.5-Omni-7B using reinforcement learning, achieving state-of-the-art performance across three types of modality input-output combinations on the ALR test set.
2 Related Work
Open-Source Datasets for Audio Reasoning.
Currently, there are very few publicly available datasets for audio reasoning. The audio datasets released in recent years typically use audio as a descriptive input, combined with text, and are fed jointly into ALMs. However, the final annotations in these datasets are purely textual (Suzgun et al., 2023; Kong et al., 2024; Ghosh et al., 2025). Recently, some researchers have begun to recognize the critical role of CoT prompting in audio reasoning datasets (Xie et al., 2025b). Nevertheless, our study is the first open-source dataset in which audio serves as the annotated modality for reasoning.
Datasets Type Transcripts CoT annotations Samples Hours
CoTA Sound, Speech, Music 1.2M 6K
AudioSkills Speech 4.2M 9.3K
LongAudio Audio 263K 8.5K
Big Bench Audio Speech 1000 2
ALR (Our) Speech 6446 1K
Table 1: Comparison of reasoning audio datasets. ALR is the only dataset providing both transcripts and chain-of-thought annotations in addition to audio, supporting step-by-step reasoning and multimodal evaluation.
Chain-of-Thought Reasoning.
LLMs enhance their reasoning ability through in-context learning (ICL), which processes prompts within the surrounding context. CoT techniques further reinforce this capability. Prominent CoT approaches include Tree-of-Thought (ToT) (Yao et al., 2023), manually crafted few-shot CoT (Wei et al., 2022), and various automatic generation strategies (Jin et al., 2024). Recent works have also examined the necessity, theoretical underpinnings, and task-specific effectiveness of CoT reasoning (Sprague et al., 2025). OpenAI’s ol model (Jaech et al., 2024) has reignited interest in CoT prompting and is often paired with reinforcement learning-based training approaches (Hu, 2025; Xie et al., 2025a).
Multimodal Chain-of-Thought Reasoning.
CoT prompting has seen notable advancements in the multimodal domain. For instance, Visual-CoT (Shao et al., 2024) integrates object detection to assist reasoning, while LLaVA-CoT (Xu et al., 2024) and MAmmoTH-VL (Guo et al., 2024) improve performance via dataset augmentation. However, CoT applications in the audio domain are still in their infancy. Audio-CoT (Ma et al., 2025) demonstrates that zero-shot CoT prompting yields improvements on simple audio tasks, but remains inadequate for complex reasoning.
Although current ALMs have made progress in comprehension and real-time response, their capability in CoT-style reasoning remains underexplored. Our study addresses this research gap by systematically investigating the application of CoT techniques within ALMs.
3 Audio Logic Reasoning Dataset
3.1 Main Features
We introduce the Audio Logic Reasoning (ALR) dataset, a novel benchmark specifically designed to advance audio-based reasoning research. Unlike most existing datasets that rely solely on textual annotations, ALR provides both audio-level and CoT annotations, enabling models to learn and reason directly from auditory inputs. The dataset focuses on speech-based scenarios, aligning with real-world applications such as spoken dialogue understanding and auditory commonsense reasoning.
As shown in Table 1, existing audio datasets such as CoTA (Xie et al., 2025b), AudioSkills (Ghosh et al., 2025), LongAudio (Ghosh et al., 2025), and Big Bench Audio (Suzgun et al., 2023) either lack audio-level annotations or do not provide CoT reasoning traces. For instance, while CoTA includes CoT-style annotations, it does not offer audio annotations/transcripts, limiting its ability to support models that operate directly on audio. Similarly, AudioSkills and LongAudio provide a large volume of audio samples but lack both reasoning annotations and step-by-step rationales.
In contrast, ALR is the only dataset in this comparison that includes both audio and CoT annotations. Although it contains fewer samples (6,446) than large-scale datasets such as AudioSkills, it is more focused on logic-oriented reasoning tasks and offers 1,074 hours of carefully annotated speech audio. This dual-modality annotation makes ALR a uniquely valuable resource for training and evaluating Logic-Aware ALMs in tasks requiring structured reasoning grounded in audio perception.
3.2 Data Generation Pipeline
Figure 2: Three-stage pipeline for constructing ALR samples. Step 1: Structured logical inputs are converted into natural prompts through colloquialization. Step 2: A large language model generates chain-of-thought reasoning and final answers. Step 3: Both input and output texts are synthesized into speech using a text-to-speech model, producing paired audio data aligned with logical reasoning.
The Audio Logic Reasoning (ALR) dataset is constructed via a structured pipeline that converts textual logical reasoning tasks into natural spoken audio interactions, enabling audio-based reasoning model development. The full process is in Figure 2.
We begin with logic problems from the LogiQA 2.0 NLI (Liu et al., 2023), which provides a structured triplet: a major premise, a minor premise, and a conclusion. These inputs are first processed through a colloquialization module to convert formal logic into natural, conversational language (e.g., Major premise is …, Let’s figure out the logical connection…). This step enhances the naturalness of the audio and simulates real user prompts.
The colloquialized content and instructions are combined into user content, which is then synthesized into speech audio using the high-quality text-to-speech model MegaTTS 3 (Jiang et al., 2025). This forms the input side of the dataset. The prompts we used can be found in Table 2.
Figure 3: An example in text from the ALR dataset. The sample consists of three components: (1) User Content, which includes a natural-language prompt and a structured logical triplet (major premise, minor premise, and conclusion); (2) Chain-of-Thought response generated by a large language model, which explains the logical reasoning behind the entailment decision; and (3) the final Answer. All components are available in both text and synthesized audio formats, supporting multimodal training and evaluation.
To generate the output, we use the large language model DeepSeek-R1 to produce CoT reasoning and final answers. These reasoning traces provide step-by-step explanations and improve supervision quality. The CoT outputs are also synthesized into audio using MegaTTS 3, producing the corresponding spoken response.
As a result, each ALR sample consists of a pair of spoken audio segments: one representing the user query and the other containing the reasoning and answer, as shown in Figure 3.
Position Prompt
System Your task is to decide if the conclusion is "entailed" or "not-entailed" based on these premises. You are a wise person who answers two-choice questions, "entailed" or "not entailed". Use plain text for thought processes and answers, not markdown or LaTeX. The thought process and response style should be colloquial, which I can then translate directly into audio using the TTS model. The final output is the Answer, nothing else, and the format is Answer: YOUR ANSWER. For example: "Answer: entailed." or "Answer: not entailed." The final answer must contain nothing else! The thought process should be very complete, careful, and cautious. When you think and generate a chain of thought, you need to test your answer from various angles.
Before the major premise Let’s figure out the logical connection between these premises and the conclusion. You have two choices: "entailed" means the conclusion must be true based on the given premises, or "not-entailed" means the conclusion can’t be true based on the premises. Here’s the setup:
Behind the conclusion Your task is to decide is the conclusion is "entailed" or "not-entailed" based on these premises.
Table 2: Instructional prompts at different positions during chain-of-thought generation. The system prompt defines the overall task, output format, and desired reasoning style for speech synthesis. Contextual prompts inserted before the major premise and after the conclusion further guide the model in interpreting logical relationships.
Split # Entailed # Not-ent. Avg. Inp. Tok. Avg. Out. Tok. Avg. Inp. Dur. (s) Avg. Out. Dur. (s)
Training 2326 2858 182 1683 62.33 608.42
Test 296 360 158 1424 57.90 586.51
Validation 264 342 155 1426 57.85 586.39
Table 3: Detailed statistics of the Audio Logic Reasoning (ALR) dataset. “# Entailed” and “# Not-ent.” indicate the number of entailed and not-entailed samples, respectively. “Avg. Inp./Out. Tok.” denotes the average input/output tokens. “Avg. Inp./Out. Dur. (s)” indicates the average duration (in seconds) of input/output sequences.
4 SoundMind Algorithm
We improve our system through iterative optimization of rule-based rewards. All $\lambda$ in the following equations are hyperparameters.
Answer Format Correctness Evaluation.
To ensure the correctness of answer formatting, we first require that the token “Answer:” must appear within the last five characters of the model response. Considering the dual-modality annotation characteristics of the ALR dataset, we design two specific format scoring methods (denoted as $S_{\text{format}}$), calculated as follows:
$\mathit{S}_{\text{format}}^{(1)}=\begin{cases}\lambda_{1},&\text{if correct for text tokens}\\ 0,&\text{otherwise},\end{cases}$ (1)
$\mathit{S}_{\text{format}}^{(2)}=\begin{cases}\lambda_{2},&\text{if correct for audio tokens}\\ 0,&\text{otherwise}.\end{cases}$ (2)
Answer Correctness Evaluation.
Once the format compliance is verified, this module evaluates the factual accuracy of the model response. Specifically, the answer score ($S_{\text{answer}}$) is computed based on the consistency between the model’s predicted answer and the ground truth answer, using the following formulation:
$\mathit{S}_{\text{answer}}=\begin{cases}\lambda_{3},&\text{if answer = ground truth}\\ 0,&\text{otherwise}.\end{cases}$ (3)
Reasoning Length Evaluation.
We additionally introduce a reward evaluation based on the reasoning length of the model response. This score ($S_{\text{len}}$) is computed by comparing the ratio of the model output length to the reference reasoning length. In light of the ALR dataset’s dual-modality annotations, we design two distinct length evaluation methods, defined as follows:
$\mathit{S}_{\text{len}}^{(1)}=\lambda_{4}\times\min\left(1,\frac{L_{\text{model}}}{L_{\text{annotation}}}\right),$ (4)
$\mathit{S}_{\text{len}}^{(2)}=\lambda_{5}\times\min\left(1,\frac{T_{\text{model}}}{T_{\text{annotation}}}\right).$ (5)
$L$ denotes the length of text tokens, and $T$ denotes the length of audio tokens.
REINFORCE++ Policy Optimization.
In our setting, the policy $\pi_{\theta}$ corresponds to a large-scale ALM, which receives an audio question as input and produces a reasoning response in either text or audio form. To optimize the model with the above composite reward, we adopt the REINFORCE++ Hu (2025)—a clipped policy-gradient method that eliminates the need for a value (critic) network while leveraging PPO-style stability and sample efficiency. Specifically, REINFORCE++ updates the policy by maximizing the following objective:
$\mathcal{J}_{\mathrm{REINFORCE++}}(\theta)=\;\mathbb{E}_{(x,y)\sim\mathcal{D},\;o_{\leq t}\sim\pi_{\theta_{\mathrm{old}}}}$ (6)
$\Bigg{[}\min\left(r_{t}(\theta)\,\hat{A}_{t},\;\operatorname{clip}\left(r_{t}(\theta),1-\epsilon,1+\epsilon\right)\,\hat{A}_{t}\right)\Bigg{]}$
where $x$ is the audio input, $y$ is the generated response, and $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\mathrm{old}}}(a_{t}|s_{t})}$ is the importance sampling ratio. $\hat{A}_{t}$ is the normalized advantage, given by: $A_{t}=R(x,y)-\beta\sum_{i=t}^{T}\mathrm{KL}(i),\hat{A}_{t}=\frac{A_{t}-\mu_{A}}{\sigma_{A}}.$ Here, $R(x,y)$ is the cumulative reward, composed of the weighted sum of the above reward terms ($S_{\text{format}}$, $S_{\text{answer}}$, $S_{\text{len}}$), and $\mathrm{KL}(i)$ denotes the token-level KL divergence between the current policy and the reference SFT model. The KL term is defined as:
$\mathrm{KL}(i)=\log\frac{\pi_{\theta}(a_{i}|s_{i})}{\pi_{\mathrm{SFT}}(a_{i}|s_{i})},$ (7)
where $\pi_{\theta}$ is the current policy being optimized, $\pi_{\mathrm{SFT}}$ is the fixed reference (supervised fine-tuned) model, $\beta$ is the KL penalty coefficient, and $\mu_{A}$, $\sigma_{A}$ are the batch mean and standard deviation for normalization. $\epsilon$ is the PPO clip parameter. REINFORCE++ combines the stability of PPO’s clipped surrogate loss with the efficiency of critic-free Monte Carlo policy gradient updates, using a token-level KL penalty and batch-normalized advantages. It trains audio-language models for end-to-end reasoning over audio questions.
5 Experiments
(a) Training set (Entailed)
(b) Test set (Entailed)
(c) Validation set (Entailed)
(d) Training set (Not-entailed)
(e) Test set (Not-entailed)
(f) Validation set (Not-entailed)
Figure 4: Histograms of audio durations across dataset splits and entailment labels. Subfigures (a–c) correspond to the training, test, and validation sets for “Entailed” samples; (d–f) show the same splits for “Not-entailed” samples.
5.1 Setup
We fine-tune the Qwen2.5-Omni-7B model on the ALR dataset using the proposed SoundMind algorithm. All experiments were conducted under a consistent hardware environment, consisting of an Intel(R) Xeon(R) Platinum 8468 CPU and 8 NVIDIA H800 GPUs, each equipped with 80 GB of memory. The weighting coefficients for the reward components were set as follows: $\lambda_{1}=1.0$, $\lambda_{2}=0.5$, $\lambda_{3}=2.0$, $\lambda_{4}=1.0$, and $\lambda_{5}=0.75$. The training procedure was executed for 50,000 steps. All other hyperparameter settings followed those used in Logic-RL Xie et al. (2025a).
To consider the dual-modality nature of both input and output in the ALR dataset, we conduct experiments across three settings: 1 Table 4: audio-only input with text-only reasoning output. We select five multimodal LLMs as baselines: MiniCPM-o (Yao et al., 2024), Gemini-Pro-V1.5 (Team et al., 2024), Baichuan-Omni-1.5 (Li et al., 2025), Qwen2-Audio (Chu et al., 2024), and Qwen2.5-Omni-7B (Xu et al., 2025). 2 Table 5: text-only input with audio-only reasoning output. Since there are few bimodal LLMs with a similar number of parameters to Qwen2.5-Omni-7B, we use Qwen2.5-Omni-7B as the baseline. 3 Table 6: audio-only input with audio-only reasoning output.
5.2 Dataset Analysis
Table 3 presents statistics of the ALR dataset, highlighting its scale, balance, and linguistic richness across training, validation, and test splits. The dataset comprises a total of 6,446 samples, with a relatively balanced distribution between entailed and not-entailed instances, ensuring fair coverage of both classes for entailment prediction.
The average length of user content is approximately 160–180 tokens across all splits, while the generated CoT responses are significantly longer, averaging over 1,400 tokens. This discrepancy underscores the dataset’s emphasis on step-by-step, explainable reasoning, which is crucial for supporting interpretability in audio-based reasoning tasks.
In terms of audio representation, the user content audio averages around one minute in duration, whereas the CoT reasoning speech spans nearly 10 minutes per sample. This extended audio format provides an excellent benchmark for evaluating models’ long-form audio understanding and reasoning capabilities.
As shown in Figure 4, the ALR dataset exhibits a wide range of duration distributions that reflect critical considerations for audio-logical reasoning research. Audio segments predominantly span 3–12 minutes (with 28.6% of training samples in the 540–720s range), balancing sufficient contextual information with manageable computational complexity. The design deliberately avoids shorter clips lacking substantive reasoning content.
The dataset demonstrates a modest class imbalance (44.9% entailed vs. 55.1% not-entailed overall), mirroring real-world reasoning scenarios. This imbalance is particularly pronounced in mid-range durations (240–420s), where audio complexity appears to challenge entailment determination. Notably, the 480–540s bin in the training set shows an inverse distribution (56.9% entailed), possibly reflecting unique acoustic features or annotation artifacts.
Strategic split allocation maintains consistent entailment ratios across training (42.8%), test (54.9%), and validation (56.3%) sets, ensuring reliable evaluation while preserving inherent complexity. Proportional representation across all duration bins in each split helps prevent duration-based bias. The near-balanced nature of the test set (360 entailed vs. 296 not-entailed) supports rigorous and fair model assessment.
Overall, this distribution profile directly supports the dataset’s objective of advancing bimodal reasoning by providing diverse yet controlled audio-text interaction scenarios. These settings challenge models to develop genuine cross-modal understanding, rather than relying on duration-specific pattern recognition.
5.3 Results and Analysis
Model Accuracy (%$\uparrow$
MiniCPM-o 73.17
Gemini-Pro-V1.5 74.54
Baichuan-Omni-1.5 70.58
Qwen2-Audio 58.23
Qwen2.5-Omni-7B 77.59
Qwen2.5-Omni-7B-RL (Our) 81.40
Table 4: Test accuracy (%) of different models on the ALR benchmark for audio-to-text logical reasoning (audio input, text output).
SoundMind achieves state-of-the-art performance on the audio-to-text reasoning task. Table 4 presents results for generating textual reasoning outputs from audio-only inputs. Our Qwen2.5-Omni-7B-RL model achieves an accuracy of 81.40%, outperforming all evaluated baselines. Compared to its supervised counterpart Qwen2.5-Omni-7B (77.59%), the reinforcement-tuned variant achieves a notable absolute gain of 3.81%. This indicates that the SoundMind reward framework significantly improves the model’s ability to generate accurate, text-based logical conclusions based on audio inputs. Among non-RL baselines, Gemini-Pro-V1.5 achieves the second-best performance with 74.54%, followed by MiniCPM-o (73.17%), and Baichuan-Omni-1.5 (70.58%). Qwen2-Audio lags behind with 58.23%, highlighting that generic multimodal alignment is insufficient for reasoning-heavy tasks without dedicated optimization. Qwen2.5-Omni-7B-RL improves upon Gemini-Pro by a margin of 6.86%, and surpasses MiniCPM-o by 8.23%, establishing a clear lead across models of comparable or larger capacity. These results validate that our reinforcement learning setup improves reasoning accuracy and generalizes effectively across audio entailment tasks, driven by structured reward signals for answer format, answer correctness, and reasoning length.
SoundMind enhances logical structuring in text-to-audio reasoning despite modality shift. In Table 5, we evaluate the model’s ability to generate speech-based reasoning from text-only prompts. The RL-enhanced model achieves 83.84% accuracy, representing a 3.05% improvement over the supervised baseline. This result highlights that even when the input lacks audio grounding, the reward structure guides the model toward more structured and consistent reasoning in speech. The increased reasoning depth, induced by chain-of-thought reinforcement signals, indicates that SoundMind is capable of refining response quality across modalities—not only in interpreting audio, but also in generating semantically coherent outputs in speech form.
Model WER (%$\downarrow$ Acc. (%$\uparrow$
Qwen2.5-Omni-7B 2.18 80.79
Qwen2.5-Omni-7B-RL (Our) 6.99 83.84
[…]
In fully speech-based settings, SoundMind generalizes reasoning without textual anchors. Table 6 investigates the most demanding configuration: generating spoken reasoning from audio-only inputs, with no access to textual content during inference. In this dual-audio scenario, our model maintains a strong performance gain, raising accuracy from 77.59% to 81.40%. This demonstrates that SoundMind can internalize abstract logical dependencies in audio alone—without relying on textual shortcuts. Such robustness is especially valuable for real-world applications involving spoken dialogue systems, voice-based tutoring agents, or accessibility tools that require full audio-audio reasoning capabilities.
Across all tasks, SoundMind consistently improves logical accuracy while maintaining acceptable speech quality. Our Qwen2.5-Omni-7B-RL model outperforms all baselines in reasoning accuracy across the three evaluated input-output modality pairs. These results confirm that reward-guided reinforcement learning can strengthen logical inference capabilities in audio-language models. While we observe a moderate rise in Word Error Rate (WER) in speech-based outputs, this trade-off is expected given the model’s increased focus on semantic correctness and structured explanation. The observed WER remains within acceptable bounds and does not hinder interpretability. Future work can explore fluency-aware RL objectives or dual-path reward balancing to further harmonize reasoning fidelity and audio naturalness.
Model WER (%$\downarrow$ Acc. (%$\uparrow$
Qwen2.5-Omni-7B 2.23 77.59
Qwen2.5-Omni-7B-RL (Our) 8.95 81.40
[…]
6 Conclusion
We introduce SoundMind, a novel rule-based reinforcement learning framework that empowers large-scale audio-language models with advanced logical reasoning capabilities across both audio and textual modalities. To enable such training, we build the Audio Logical Reasoning (ALR) dataset, a dual-modality benchmark comprising 6,446 high-quality samples annotated with chain-of-thought reasoning in both audio and text forms. Experimental results demonstrate that our method significantly improves performance and establishes new state-of-the-art results on the ALR benchmark, consistently outperforming strong baselines across three reasoning settings: text-to-audio, audio-to-text, and audio-to-audio. We hope this work can serve as a foundation for future research in logic-oriented large-scale audio-language modeling and in reinforcement-driven multimodal reasoning.
Limitations
While our approach demonstrates strong performance, certain aspects warrant further exploration. In particular, the Word Error Rate (WER) in audio generation remains competitive, but still leaves room for improving reasoning fidelity during speech synthesis. Addressing this limitation could lead to more robust, adaptive, and human-aligned audio reasoning systems in future work.
Ethical Considerations
We have not identified any ethical concerns directly related to this study.
Acknowledgment
This study is supported by the Department of Defense grant HT9425-23-1-0267.

### v2

@@
 Abstract
-While large language models have shown reasoning capabilities, their application to the audio modality, particularly in large audio-language models (ALMs), remains significantly underdeveloped.
-Addressing this gap requires a systematic approach, involving a capable base model, high-quality reasoning-oriented audio data, and effective training algorithms.
-In this study, we present a comprehensive solution: we introduce the Audio Logical Reasoning (ALR) dataset, consisting of 6,446 text-audio annotated samples specifically designed for complex reasoning tasks.
-Building on this resource, we propose SoundMind, a rule-based reinforcement learning (RL) algorithm tailored to endow ALMs with deep bimodal reasoning abilities.
-By training Qwen2.5-Omni-7B on the ALR dataset using SoundMind, our approach achieves state-of-the-art performance in audio logical reasoning.
-This work highlights the impact of combining high-quality, reasoning-focused datasets with specialized RL techniques, advancing the frontier of auditory intelligence in language models.
-Our code and the proposed dataset are available at https://github.com/xid32/SoundMind.
+While large language models have demonstrated impressive reasoning abilities, their extension to the audio modality, particularly within large audio-language models (LALMs), remains underexplored.
+Addressing this gap requires a systematic approach that involves a capable base model, high-quality reasoning-oriented audio data, and effective training algorithms.
+In this work, we present a comprehensive solution for audio logical reasoning (ALR) tasks: we introduce SoundMind, a dataset of 6,446 audio–text annotated samples specifically curated to support complex reasoning.
+Building on this resource, we propose SoundMind-RL, a rule-based reinforcement learning (RL) algorithm designed to equip audio-language models with robust audio–text reasoning capabilities.
+By fine-tuning Qwen2.5-Omni-7B on the proposed SoundMind dataset using SoundMind-RL, we achieve strong and consistent improvements over state-of-the-art baselines on the SoundMind benchmark.
+This work highlights the benefit of combining high-quality, reasoning-focused datasets with specialized RL techniques, and contributes to advancing auditory intelligence in language models.
+The code and dataset are publicly available at https://github.com/xid32/SoundMind.
 SoundMind: RL-Incentivized Logic Reasoning for Audio-Language Models
-Xingjian Diao1, Chunhui Zhang1, Keyi Kong2, Weiyi Wu1, Chiyu Ma1, Zhongyu Ouyang1, Peijun Qing1, Soroush Vosoughi1, Jiang Gui1 1Dartmouth College, 2Shandong University xingjian.diao.gr@dartmouth.edu
+Xingjian Diao, Chunhui Zhang, Keyi Kong , Weiyi Wu, Chiyu Ma, Zhongyu Ouyang, Peijun Qing, Soroush Vosoughi, Jiang Gui Dartmouth College, Shandong University xingjian.diao.gr@dartmouth.edu [Uncaptioned image]SoundMind Dataset
 1 Introduction
-Figure 1: Overview of the end-to-end bimodal Natural Language Inference (NLI) task enabled by large audio-language models.
-The model receives input and produces output in either pure text or pure audio format.
+Figure 1: Overview of a SoundMind sample for the audio logical reasoning (ALR) task.
+Each instance contains User Content (top), including a natural-language prompt, structured logical triplets, and a question, as well as Response (bottom), which consists of step-by-step chain-of-thought reasoning and the final answer.
+Both components are provided in text and accompanied by the corresponding input–output speech audio (right).
+SoundMind therefore offers complete logical reasoning tasks with fully aligned content and response across text and audio modalities, supporting multimodal training and evaluation of large audio–language models.
 Large language models (LLMs) have made remarkable strides in reasoning capabilities through innovations such as Chain-of-Thought (CoT) prompting and specialized reasoning architectures.
-Recent models such as OpenAI’s o1 model (Jaech et al., 2024) and Deepseek-R1 (Guo et al., 2025) have demonstrated exceptional performance on complex mathematical and programming tasks (Zhao et al., 2024; Team et al., 2025; Muennighoff et al., 2025).
+Recent models such as OpenAI’s o1 model (Jaech et al., 2024) and Deepseek-R1 (Guo et al., 2025) have demonstrated exceptional performance on complex logical tasks (Kimi Team et al., 2025; Zhao et al., 2024; Yuan et al., 2025; Zhang et al., 2024, 2025a, 2025b; Han et al., 2024).
 Particularly noteworthy is Deepseek-R1’s rule-based reinforcement learning approach, which enables emergent reasoning without relying on traditional frameworks like Monte Carlo Tree Search (Wan et al., 2024) or process reward models (Lightman et al., 2023).
-This general reasoning paradigm has successfully extended to the visual domain, where frameworks like Visual-CoT (Shao et al., 2024) have significantly enhanced multimodal models’ cognitive abilities for image and video reasoning.
-Despite these advances in text and visual reasoning, a significant gap exists in audio reasoning and generation capabilities.
-While audio-language models (ALMs) like Audio Flamingo (Kong et al., 2024) and Qwen2.5-Omni-7B (Xu et al., 2025) have made progress in audio understanding, end-to-end audio reasoning remains underdeveloped.
+This general reasoning paradigm has been successfully extended to the visual domain, where frameworks such as Visual-CoT (Shao et al., 2024) have significantly improved multimodal models’ cognitive abilities for image and video reasoning.
+Despite these advances in text and visual reasoning, there is a significant gap in audio reasoning and generation capabilities.
+While large audio-language models (LALMs) like Audio Flamingo (Kong et al., 2024) and Qwen2.5-Omni-7B (Xu et al., 2025) have made progress in audio understanding, end-to-end audio reasoning remains underdeveloped.
 This limitation stems primarily from two factors: (1) the simplicity of existing audio reasoning datasets (Suzgun et al., 2023; Kong et al., 2024), which often contain only brief textual labels without proper audio modality annotations, and (2) the technical challenges in maintaining reasoning coherence during long-duration audio generation.
 Current CoT methods applied to audio often lead to hallucinations and performance degradation when generating extended reasoning sequences.
-The lack of aligned audio-text annotations further impedes research on reasoning-driven audio generation, creating a bottleneck in developing ALMs with sophisticated reasoning capabilities.
-To address these challenges in audio reasoning (Figure 1), we introduce a comprehensive approach with two key components.
-First, we construct the Audio Logical Reasoning (ALR) dataset (Liu et al., 2023), containing 6,446 dual-modality samples with user content, CoT reasoning steps, final answers, and corresponding audio input/output.
-Derived from the Logi-QA 2.0-NLI dataset, ALR provides complete logical reasoning tasks with both content and response in text and audio modalities.
-Second, we propose the SoundMind algorithm, which builds upon the ALR dataset to address performance issues in long-form reasoning audio generation.
-Drawing inspiration from the Logic-RL framework (Xie et al., 2025a), SoundMind incorporates a strict format-based reward mechanism to prevent shortcut reasoning biased toward the textual modality.
-It leverages the REINFORCE++ algorithm (Hu, 2025) alongside the reward design principles from Deepseek-R1 for effective post-training.
-Our main contributions are as follows:
-  • ALR Dataset: We release a high-quality dual-modality dataset containing 6,446 samples with user content, CoT reasoning, answers, and corresponding audio input/output.
-The ALR dataset includes deep reasoning annotations in both text and audio formats, providing a valuable resource for RL-based training of audio-language models.
-  • SoundMind RL Algorithm: Taking into account the unique challenges of audio reasoning generation, we design a strict format-based reward mechanism that maintains reasoning coherence across modalities.
-By combining the improved REINFORCE++ algorithm with the reward design principles of Deepseek-R1, we fine-tune Qwen2.5-Omni-7B using reinforcement learning, achieving state-of-the-art performance across three types of modality input-output combinations on the ALR test set.
-2 Related Work
-Open-Source Datasets for Audio Reasoning.
-Currently, there are very few publicly available datasets for audio reasoning.
-The audio datasets released in recent years typically use audio as a descriptive input, combined with text, and are fed jointly into ALMs.
-However, the final annotations in these datasets are purely textual (Suzgun et al., 2023; Kong et al., 2024; Ghosh et al., 2025).
-Recently, some researchers have begun to recognize the critical role of CoT prompting in audio reasoning datasets (Xie et al., 2025b).
-Nevertheless, our study is the first open-source dataset in which audio serves as the annotated modality for reasoning.
+The lack of aligned audio-text annotations further impedes research on reasoning-driven audio generation, creating a bottleneck in developing LALMs with sophisticated reasoning capabilities.
+To address these challenges in audio reasoning, we introduce a comprehensive approach with two key components.
+First, we introduce SoundMind, a benchmark dataset specifically designed for the audio logical reasoning (ALR) task.
+It contains 6,446 audio–text aligned samples, each containing user content, step-by-step chain-of-thought reasoning, final answers, and the corresponding input–output speech audio (Figure 1).
+Built upon the LogiQA 2.0–NLI dataset (Liu et al., 2023), SoundMind preserves the full logical structure and augments it with parallel text–audio annotations, enabling models to learn reasoning grounded in audio.
+Second, we propose SoundMind-RL, a rule-based reinforcement learning algorithm that addresses the challenges of long-form audio reasoning generation.
+Drawing inspiration from the Logic-RL framework (Xie et al., 2025a), SoundMind-RL incorporates a strict format-based reward mechanism to prevent shortcut reasoning biased toward the textual modality.
+It leverages the REINFORCE++ algorithm (Hu et al., 2025) alongside the reward design principles from Deepseek-R1 for effective post-training.
+Our main contributions are summarized as follows:
+  • SoundMind Dataset: We release a high-quality audio logical reasoning dataset comprising 6,446 samples with user content, CoT reasoning, answers and corresponding audio input/output.
+It provides deep reasoning annotations in both text and audio modalities, serving as a valuable resource for RL-based training and evaluation of audio–language models.
+  • SoundMind-RL Algorithm: Taking into account the unique challenges of audio reasoning generation, we design a strict format-based reward mechanism that maintains reasoning coherence across modalities.
+By combining the improved REINFORCE++ algorithm with reward design principles from DeepSeek-R1, we fine-tune Qwen2.5-Omni-7B through reinforcement learning, achieving state-of-the-art results across three input–output modality configurations on the SoundMind benchmark.
 Datasets Type Transcripts CoT annotations Samples Hours
@@
 Big Bench Audio Speech 1000 2
-ALR (Our) Speech 6446 1K
-Table 1: Comparison of reasoning audio datasets.
-ALR is the only dataset providing both transcripts and chain-of-thought annotations in addition to audio, supporting step-by-step reasoning and multimodal evaluation.
+SoundMind (Ours) Speech 6446 1K
+Table 1: Comparison of publicly available datasets for audio reasoning.
+Existing resources differ in scope: some provide speech data without transcripts, while others include transcripts but lack step-by-step reasoning annotations aligned with audio.
+These gaps make systematic study of multimodal reasoning challenging.
+SoundMind combines transcripts, chain-of-thought (CoT) annotations, and speech audio into a unified benchmark, offering a balanced resource for training and evaluating audio–language models.
+2 Related Work
+Open-Source Datasets for Audio Reasoning.
+Research on audio reasoning remains limited by the scarcity of publicly available benchmarks, making systematic evaluation and comparison difficult.
+Existing corpora typically treat audio as descriptive input paired with text, yet their supervision remains purely textual, providing final annotations exclusively in text form (Suzgun et al., 2023; Kong et al., 2024; Ghosh et al., 2025).
+More recently, a number of studies have begun to recognize the critical role of chain-of-thought (CoT) prompting in audio reasoning datasets (Xie et al., 2025b), highlighting the importance of structured reasoning traces.
+To the best of our knowledge, SoundMind is among the first open-source resources where audio itself serves as the annotated modality for reasoning, providing parallel text–audio representations and enabling models to learn from and be evaluated on spoken reasoning traces.
 Chain-of-Thought Reasoning.
@@
 Recent works have also examined the necessity, theoretical underpinnings, and task-specific effectiveness of CoT reasoning (Sprague et al., 2025).
-OpenAI’s ol model (Jaech et al., 2024) has reignited interest in CoT prompting and is often paired with reinforcement learning-based training approaches (Hu, 2025; Xie et al., 2025a).
+OpenAI’s ol model (Jaech et al., 2024) has reignited interest in CoT prompting and is often paired with reinforcement learning-based training approaches (Hu et al., 2025; Xie et al., 2025a).
 Multimodal Chain-of-Thought Reasoning.
@@
 Audio-CoT (Ma et al., 2025) demonstrates that zero-shot CoT prompting yields improvements on simple audio tasks, but remains inadequate for complex reasoning.
-Although current ALMs have made progress in comprehension and real-time response, their capability in CoT-style reasoning remains underexplored.
+Although current audio-language models (ALMs) have made progress in comprehension and real-time response, their capability in CoT-style reasoning remains underexplored.
 Our study addresses this research gap by systematically investigating the application of CoT techniques within ALMs.
-3 Audio Logic Reasoning Dataset
-3.1 Main Features
-We introduce the Audio Logic Reasoning (ALR) dataset, a novel benchmark specifically designed to advance audio-based reasoning research.
-Unlike most existing datasets that rely solely on textual annotations, ALR provides both audio-level and CoT annotations, enabling models to learn and reason directly from auditory inputs.
-The dataset focuses on speech-based scenarios, aligning with real-world applications such as spoken dialogue understanding and auditory commonsense reasoning.
-As shown in Table 1, existing audio datasets such as CoTA (Xie et al., 2025b), AudioSkills (Ghosh et al., 2025), LongAudio (Ghosh et al., 2025), and Big Bench Audio (Suzgun et al., 2023) either lack audio-level annotations or do not provide CoT reasoning traces.
-For instance, while CoTA includes CoT-style annotations, it does not offer audio annotations/transcripts, limiting its ability to support models that operate directly on audio.
-Similarly, AudioSkills and LongAudio provide a large volume of audio samples but lack both reasoning annotations and step-by-step rationales.
-In contrast, ALR is the only dataset in this comparison that includes both audio and CoT annotations.
-Although it contains fewer samples (6,446) than large-scale datasets such as AudioSkills, it is more focused on logic-oriented reasoning tasks and offers 1,074 hours of carefully annotated speech audio.
-This dual-modality annotation makes ALR a uniquely valuable resource for training and evaluating Logic-Aware ALMs in tasks requiring structured reasoning grounded in audio perception.
+3 Dataset Design and Construction
+3.1 Key Characteristics
+We introduce SoundMind, a unified benchmark for the audio logical reasoning (ALR) task, specifically developed to advance research on reasoning grounded in auditory input.
+In contrast to most existing datasets that provide only textual supervision, SoundMind offers both text-level and audio-level chain-of-thought (CoT) annotations, allowing models to learn and reason directly from spoken prompts and responses.
+The dataset emphasizes speech-based scenarios, making it well suited for applications such as dialogue understanding and auditory commonsense reasoning.
+As shown in Table 1, existing audio datasets such as CoTA (Xie et al., 2025b), AudioSkills (Ghosh et al., 2025), LongAudio (Ghosh et al., 2025), and Big Bench Audio (Suzgun et al., 2023) provide valuable resources but remain limited in their support for auditory reasoning.
+Some lack audio-level annotations entirely, while others include audio but omit chain-of-thought (CoT) reasoning traces or detailed step-by-step rationales.
+For instance, although CoTA offers CoT-style annotations, it does not include audio transcripts, which limits its utility for training and evaluating models that operate directly on speech input.
+Likewise, AudioSkills and LongAudio supply a large quantity of audio data but no reasoning supervision, making them less suitable for investigating step-by-step inference or developing models that require explicit alignment between input and reasoning output.
+Figure 2: Three-step pipeline for constructing SoundMind samples.
+Step 1: Structured logical triplets are converted into natural, conversational prompts through a Colloquialization module.
+Step 2: A large language model generates detailed chain-of-thought reasoning and final answers, providing rich supervision for reasoning tasks.
+Step 3: Both the user content and the generated reasoning–answer pairs are converted into speech using a text-to-speech (TTS) model, yielding two carefully aligned audio segments that together capture the full reasoning interaction.
+The SoundMind dataset is constructed through a structured pipeline that transforms textual logical reasoning tasks into natural spoken audio interactions, thereby enabling the development of audio-based reasoning models.
+The complete pipeline is illustrated in Figure 2.
+Although SoundMind contains fewer samples (6,446) than large-scale datasets such as AudioSkills, it is specifically designed for logic-oriented reasoning and provides over 1,074 hours of carefully annotated speech audio with both text-level and audio-level CoT supervision.
+This comprehensive annotation makes SoundMind a valuable benchmark for training and evaluating models that require structured, multimodal reasoning grounded in audio perception.
 3.2 Data Generation Pipeline
-Figure 2: Three-stage pipeline for constructing ALR samples.
-Step 1: Structured logical inputs are converted into natural prompts through colloquialization.
-Step 2: A large language model generates chain-of-thought reasoning and final answers.
-Step 3: Both input and output texts are synthesized into speech using a text-to-speech model, producing paired audio data aligned with logical reasoning.
-The Audio Logic Reasoning (ALR) dataset is constructed via a structured pipeline that converts textual logical reasoning tasks into natural spoken audio interactions, enabling audio-based reasoning model development.
-The full process is in Figure 2.
-We begin with logic problems from the LogiQA 2.0 NLI (Liu et al., 2023), which provides a structured triplet: a major premise, a minor premise, and a conclusion.
-These inputs are first processed through a colloquialization module to convert formal logic into natural, conversational language (e.g., Major premise is …, Let’s figure out the logical connection…).
-This step enhances the naturalness of the audio and simulates real user prompts.
-The colloquialized content and instructions are combined into user content, which is then synthesized into speech audio using the high-quality text-to-speech model MegaTTS 3 (Jiang et al., 2025).
-This forms the input side of the dataset.
-The prompts we used can be found in Table 2.
-Figure 3: An example in text from the ALR dataset.
-The sample consists of three components: (1) User Content, which includes a natural-language prompt and a structured logical triplet (major premise, minor premise, and conclusion); (2) Chain-of-Thought response generated by a large language model, which explains the logical reasoning behind the entailment decision; and (3) the final Answer.
-All components are available in both text and synthesized audio formats, supporting multimodal training and evaluation.
-To generate the output, we use the large language model DeepSeek-R1 to produce CoT reasoning and final answers.
-These reasoning traces provide step-by-step explanations and improve supervision quality.
-The CoT outputs are also synthesized into audio using MegaTTS 3, producing the corresponding spoken response.
-As a result, each ALR sample consists of a pair of spoken audio segments: one representing the user query and the other containing the reasoning and answer, as shown in Figure 3.
+Our pipeline takes as input structured logical triplets consisting of a major premise, a minor premise, and a conclusion.
+These are first processed by a Colloquialization module that rewrites the formal statements into natural, conversational prompts (e.g., “The major premise is …”, “Let’s figure out the logical connection …”).
+This step enhances the fluency of the resulting audio and more closely reflects realistic user queries.
 Position Prompt
@@
 Behind the conclusion Your task is to decide is the conclusion is "entailed" or "not-entailed" based on these premises.
-Table 2: Instructional prompts at different positions during chain-of-thought generation.
-The system prompt defines the overall task, output format, and desired reasoning style for speech synthesis.
-Contextual prompts inserted before the major premise and after the conclusion further guide the model in interpreting logical relationships.
-Split # Entailed # Not-ent. Avg. Inp. Tok. Avg. Out. Tok. Avg. Inp. Dur. (s) Avg. Out. Dur. (s)
-Training 2326 2858 182 1683 62.33 608.42
-Test 296 360 158 1424 57.90 586.51
-Validation 264 342 155 1426 57.85 586.39
-Table 3: Detailed statistics of the Audio Logic Reasoning (ALR) dataset.
-“# Entailed” and “# Not-ent.” indicate the number of entailed and not-entailed samples, respectively.
-“Avg.
-Inp./Out.
-Tok.” denotes the average input/output tokens.
-“Avg.
-Inp./Out.
-Dur.
-(s)” indicates the average duration (in seconds) of input/output sequences.
-4 SoundMind Algorithm
+Table 2: Instructional prompts used at different positions during chain-of-thought generation.
+The system prompt specifies the overall task, output format, and encourages a careful, colloquial reasoning style suitable for speech synthesis.
+Additional contextual prompts placed before the major premise and after the conclusion provide further guidance, helping the model interpret logical relations and produce well-structured reasoning traces.
+Figure 3: An illustrative example from the SoundMind dataset.
+Each sample contains three components: (1) User Content, which includes a natural-language prompt and a structured logical triplet (major premise, minor premise, conclusion); (2) Chain-of-Thought, a step-by-step reasoning trace generated by a large language model, explaining the entailment decision; and (3) the final Answer.
+All components are provided in both text form and synthesized speech audio, enabling fully multimodal training and evaluation of audio–language models.
+The colloquialized content and instructions are merged into a single User Content block and synthesized into speech audio using the high-fidelity text-to-speech model MegaTTS 3 (Jiang et al., 2025), forming the input side of the dataset.
+The prompts used in this process are listed in Table 2.
+To generate the output, we employ the large language model DeepSeek-R1 to produce step-by-step CoT reasoning and final answers, providing richer supervision for model training.
+The generated reasoning and answers are then synthesized into speech audio with MegaTTS 3 (Jiang et al., 2025), yielding the corresponding spoken response.
+As a result, each SoundMind instance consists of two carefully aligned audio segments: one representing the user query and the other presenting the detailed step-by-step reasoning and the corresponding final answer, as illustrated in Figure 3.
+4 SoundMind-RL Algorithm
 We improve our system through iterative optimization of rule-based rewards. All $\lambda$ in the following equations are hyperparameters.
@@
 To ensure the correctness of answer formatting, we first require that the token “Answer:” must appear within the last five characters of the model response.
-Considering the dual-modality annotation characteristics of the ALR dataset, we design two specific format scoring methods (denoted as $S_{\text{format}}$), calculated as follows:
+Considering that the SoundMind dataset covers both text and audio modalities, we design two specific format scoring methods (denoted as $S_{\text{format}}$), calculated as follows:
 $\mathit{S}_{\text{format}}^{(1)}=\begin{cases}\lambda_{1},&\text{if correct for text tokens}\\ 0,&\text{otherwise},\end{cases}$ (1)
@@
 This score ($S_{\text{len}}$) is computed by comparing the ratio of the model output length to the reference reasoning length.
-In light of the ALR dataset’s dual-modality annotations, we design two distinct length evaluation methods, defined as follows:
+Given that the SoundMind dataset provides supervision in both text and audio modalities, we design two distinct length evaluation methods, defined as follows:
 $\mathit{S}_{\text{len}}^{(1)}=\lambda_{4}\times\min\left(1,\frac{L_{\text{model}}}{L_{\text{annotation}}}\right),$ (4)
 $\mathit{S}_{\text{len}}^{(2)}=\lambda_{5}\times\min\left(1,\frac{T_{\text{model}}}{T_{\text{annotation}}}\right).$ (5)
-$L$ denotes the length of text tokens, and $T$ denotes the length of audio tokens.
+where $L$ denotes the length of text tokens, and $T$ denotes the length of audio tokens.
 REINFORCE++ Policy Optimization.
 In our setting, the policy $\pi_{\theta}$ corresponds to a large-scale ALM, which receives an audio question as input and produces a reasoning response in either text or audio form.
-To optimize the model with the above composite reward, we adopt the REINFORCE++ Hu (2025)—a clipped policy-gradient method that eliminates the need for a value (critic) network while leveraging PPO-style stability and sample efficiency.
+To optimize the model with the above composite reward, we adopt the REINFORCE++ Hu et al.
+(2025)—a clipped policy-gradient method that eliminates the need for a value (critic) network while leveraging PPO-style stability and sample efficiency.
 Specifically, REINFORCE++ updates the policy by maximizing the following objective:
@@
 5 Experiments
+Split # Entailed # Not-ent. Avg. Inp. Tok. Avg. Out. Tok. Avg. Inp. Dur. (s) Avg. Out. Dur. (s)
+Training 2326 2858 182 1683 62.33 608.42
+Test 296 360 158 1424 57.90 586.51
+Validation 264 342 155 1426 57.85 586.39
+Table 3: Detailed statistics of the SoundMind dataset across training, validation, and test splits.
+“# Entailed” and “# Not-ent.” denote the number of samples labeled as entailed and not-entailed.
+“Avg.
+Inp./Out.
+Tok.” reports the mean token counts for the input and outputs, respectively.
+“Avg.
+Inp./Out.
+Dur.
+(s)” provides the average duration (in seconds) of the input audio and the corresponding output audio.
+Together, these statistics highlight the dataset’s scale, near-balanced class distribution, and focus on long-form reasoning supervision.
 (a) Training set (Entailed)
@@
 (f) Validation set (Not-entailed)
-Figure 4: Histograms of audio durations across dataset splits and entailment labels.
-Subfigures (a–c) correspond to the training, test, and validation sets for “Entailed” samples; (d–f) show the same splits for “Not-entailed” samples.
+Figure 4: Distribution of audio durations across dataset splits and entailment labels.
+Subfigures (a–c) show histograms for “Entailed” samples in the training, test, and validation sets, while (d–f) show the corresponding distributions for “Not-entailed” samples.
+These plots illustrate the broad coverage of audio durations and the consistent duration profiles across splits, highlighting that the SoundMind dataset offers both diversity and balance for evaluating model performance across different temporal contexts.
 5.1 Setup
-We fine-tune the Qwen2.5-Omni-7B model on the ALR dataset using the proposed SoundMind algorithm.
+We fine-tune the Qwen2.5-Omni-7B model on the SoundMind dataset using the proposed SoundMind-RL algorithm.
 All experiments were conducted under a consistent hardware environment, consisting of an Intel(R) Xeon(R) Platinum 8468 CPU and 8 NVIDIA H800 GPUs, each equipped with 80 GB of memory.
@@
 (2025a).
-To consider the dual-modality nature of both input and output in the ALR dataset, we conduct experiments across three settings: 1 Table 4: audio-only input with text-only reasoning output.
-We select five multimodal LLMs as baselines: MiniCPM-o (Yao et al., 2024), Gemini-Pro-V1.5 (Team et al., 2024), Baichuan-Omni-1.5 (Li et al., 2025), Qwen2-Audio (Chu et al., 2024), and Qwen2.5-Omni-7B (Xu et al., 2025).
-2 Table 5: text-only input with audio-only reasoning output.
-Since there are few bimodal LLMs with a similar number of parameters to Qwen2.5-Omni-7B, we use Qwen2.5-Omni-7B as the baseline.
-3 Table 6: audio-only input with audio-only reasoning output.
+To fully exploit the availability of both text and audio modalities in the SoundMind dataset, we evaluate our approach under three input–output configurations: 1 Table 4: audio-only input with text-based reasoning output, benchmarked against five strong multimodal LLMs: MiniCPM-o (Yao et al., 2024), Gemini-Pro-V1.5 (Team et al., 2024), Baichuan-Omni-1.5 (Li et al., 2025), Qwen2-Audio (Chu et al., 2024), and Qwen2.5-Omni-7B (Xu et al., 2025).
+2 Table 5: text-only input with audio-based reasoning output.
+Due to the lack of comparable open-source multimodal models, we use Qwen2.5-Omni-7B as the reference baseline.
+3 Table 6: audio-only input with audio-based reasoning output.
+Together, these three configurations allow us to comprehensively assess SoundMind-RL across text and audio modalities, covering fully cross-modal reasoning scenarios.
 5.2 Dataset Analysis
-Table 3 presents statistics of the ALR dataset, highlighting its scale, balance, and linguistic richness across training, validation, and test splits.
-The dataset comprises a total of 6,446 samples, with a relatively balanced distribution between entailed and not-entailed instances, ensuring fair coverage of both classes for entailment prediction.
-The average length of user content is approximately 160–180 tokens across all splits, while the generated CoT responses are significantly longer, averaging over 1,400 tokens.
-This discrepancy underscores the dataset’s emphasis on step-by-step, explainable reasoning, which is crucial for supporting interpretability in audio-based reasoning tasks.
-In terms of audio representation, the user content audio averages around one minute in duration, whereas the CoT reasoning speech spans nearly 10 minutes per sample.
-This extended audio format provides an excellent benchmark for evaluating models’ long-form audio understanding and reasoning capabilities.
-As shown in Figure 4, the ALR dataset exhibits a wide range of duration distributions that reflect critical considerations for audio-logical reasoning research.
-Audio segments predominantly span 3–12 minutes (with 28.6% of training samples in the 540–720s range), balancing sufficient contextual information with manageable computational complexity.
-The design deliberately avoids shorter clips lacking substantive reasoning content.
-The dataset demonstrates a modest class imbalance (44.9% entailed vs.
-55.1% not-entailed overall), mirroring real-world reasoning scenarios.
-This imbalance is particularly pronounced in mid-range durations (240–420s), where audio complexity appears to challenge entailment determination.
-Notably, the 480–540s bin in the training set shows an inverse distribution (56.9% entailed), possibly reflecting unique acoustic features or annotation artifacts.
-Strategic split allocation maintains consistent entailment ratios across training (42.8%), test (54.9%), and validation (56.3%) sets, ensuring reliable evaluation while preserving inherent complexity.
-Proportional representation across all duration bins in each split helps prevent duration-based bias.
-The near-balanced nature of the test set (360 entailed vs.
-296 not-entailed) supports rigorous and fair model assessment.
-Overall, this distribution profile directly supports the dataset’s objective of advancing bimodal reasoning by providing diverse yet controlled audio-text interaction scenarios.
-These settings challenge models to develop genuine cross-modal understanding, rather than relying on duration-specific pattern recognition.
+Table 3 summarizes key statistics of the SoundMind dataset, clearly illustrating its overall scale, class balance, and linguistic diversity across training, validation, and test splits.
+The corpus contains 6,446 samples with a near-balanced distribution of entailed and not-entailed instances, thereby ensuring representative coverage of both classes for robust entailment prediction.
+The user content averages 160–180 tokens across all splits, whereas the generated CoT responses are much longer, exceeding 1,400 tokens on average.
+This substantial length gap reflects the dataset’s focus on detailed, step-by-step reasoning, providing rich supervision that encourages faithful and transparent reasoning in audio-based tasks.
+In terms of audio representation, the user content speech averages roughly one minute per sample, whereas the CoT reasoning audio extends to nearly ten minutes on average.
+This substantial duration provides a challenging and realistic benchmark for assessing models’ ability to perform long-form audio comprehension and sustained logical reasoning.
+As shown in Figure 4, the SoundMind dataset covers a broad range of audio durations, capturing important variability for developing and evaluating models for audio-based logical reasoning.
+Most segments fall within 3–12 minutes, with 28.6% of training samples concentrated in the 540–720s range, which provides ample contextual information while keeping computational cost tractable.
+Notably, the dataset deliberately excludes very short clips to ensure that each sample contains sufficiently meaningful reasoning content.
+The dataset exhibits a class distribution of 44.9% entailed and 55.1% not-entailed instances overall, closely resembling patterns in many real-world reasoning scenarios.
+This skew becomes more pronounced in the mid-range durations (240–420s), suggesting that longer and acoustically richer segments may introduce additional difficulty for reliable entailment classification.
+Notably, the 480–540s range in the training split shows a reversed distribution, with 56.9% entailed samples, potentially reflecting distinctive acoustic patterns or systematic properties of the original source data.
+The SoundMind dataset is split to maintain consistent entailment ratios across training (42.8%), test (54.9%), and validation (56.3%) sets, thereby enabling reliable evaluation while preserving the inherent task complexity.
+Each split carefully preserves proportional coverage across all duration bins, effectively minimizing potential duration-related bias and ensuring representative diversity.
+The test set is nearly balanced (360 entailed vs.
+296 not-entailed), providing a solid basis for rigorous and fair model assessment.
+Overall, this distribution profile aligns with the dataset’s goal of advancing audio–language reasoning research by offering diverse but systematically controlled interaction scenarios.
+Such a design compels models to acquire robust cross-modal reasoning skills, discouraging reliance on superficial duration-based correlations and encouraging deeper semantic understanding.
 5.3 Results and Analysis
-Model Accuracy (%$\uparrow$
+5.3.1 Audio-to-Text Reasoning
+[…]
 MiniCPM-o 73.17
@@
 Qwen2.5-Omni-7B 77.59
-Qwen2.5-Omni-7B-RL (Our) 81.40
-Table 4: Test accuracy (%) of different models on the ALR benchmark for audio-to-text logical reasoning (audio input, text output).
-SoundMind achieves state-of-the-art performance on the audio-to-text reasoning task.
-Table 4 presents results for generating textual reasoning outputs from audio-only inputs.
-Our Qwen2.5-Omni-7B-RL model achieves an accuracy of 81.40%, outperforming all evaluated baselines.
-Compared to its supervised counterpart Qwen2.5-Omni-7B (77.59%), the reinforcement-tuned variant achieves a notable absolute gain of 3.81%.
-This indicates that the SoundMind reward framework significantly improves the model’s ability to generate accurate, text-based logical conclusions based on audio inputs.
-Among non-RL baselines, Gemini-Pro-V1.5 achieves the second-best performance with 74.54%, followed by MiniCPM-o (73.17%), and Baichuan-Omni-1.5 (70.58%).
-Qwen2-Audio lags behind with 58.23%, highlighting that generic multimodal alignment is insufficient for reasoning-heavy tasks without dedicated optimization.
-Qwen2.5-Omni-7B-RL improves upon Gemini-Pro by a margin of 6.86%, and surpasses MiniCPM-o by 8.23%, establishing a clear lead across models of comparable or larger capacity.
-These results validate that our reinforcement learning setup improves reasoning accuracy and generalizes effectively across audio entailment tasks, driven by structured reward signals for answer format, answer correctness, and reasoning length.
-SoundMind enhances logical structuring in text-to-audio reasoning despite modality shift.
-In Table 5, we evaluate the model’s ability to generate speech-based reasoning from text-only prompts.
-The RL-enhanced model achieves 83.84% accuracy, representing a 3.05% improvement over the supervised baseline.
-This result highlights that even when the input lacks audio grounding, the reward structure guides the model toward more structured and consistent reasoning in speech.
-The increased reasoning depth, induced by chain-of-thought reinforcement signals, indicates that SoundMind is capable of refining response quality across modalities—not only in interpreting audio, but also in generating semantically coherent outputs in speech form.
-Model WER (%$\downarrow$ Acc. (%$\uparrow$
+Qwen2.5-Omni-7B (SoundMind-RL) 81.40
+Table 4: Accuracy (%) of evaluated models on the SoundMind benchmark under the audio-to-text configuration.
+In this setting, models receive audio-only inputs and produce textual outputs containing the final answer.
+The table compares several strong multimodal LLM baselines with the reinforcement-tuned Qwen2.5-Omni-7B (SoundMind-RL).
+Qwen2.5-Omni-7B fine-tuned with SoundMind-RL achieves state-of-the-art performance on the challenging audio-to-text reasoning task.
+Table 4 reports results for generating textual outputs from audio-only inputs.
+Our model attains an accuracy of 81.40%, clearly outperforming all evaluated baselines.
+Compared to its counterpart Qwen2.5-Omni-7B (77.59%), the reinforcement-tuned variant yields an absolute improvement of 3.81%, clearly demonstrating that the SoundMind reward framework substantially enhances the model’s ability to produce accurate text-based logical conclusions from speech audio input.
+Among the remaining baseline models, Gemini-Pro-V1.5 reaches 74.54%, followed closely by MiniCPM-o (73.17%) and Baichuan-Omni-1.5 (70.58%).
+Qwen2-Audio performs the weakest, achieving only 58.23%, clearly underscoring that generic multimodal alignment is insufficient for reasoning-intensive tasks without specialized optimization.
+Overall, Qwen2.5-Omni-7B (SoundMind-RL) further improves upon Gemini-Pro by 6.86% and surpasses MiniCPM-o by 8.23%, establishing a consistently clear lead across models of comparable or larger capacity.
+These findings strongly confirm that our reinforcement learning approach strengthens reasoning accuracy and generalizes effectively across audio reasoning tasks.
+5.3.2 Text-to-Audio Reasoning
+[…]
 Qwen2.5-Omni-7B 2.18 80.79
-Qwen2.5-Omni-7B-RL (Our) 6.99 83.84
+Qwen2.5-Omni-7B (SoundMind-RL) 6.99 83.84
+Table 5: Performance of Qwen2.5-Omni-7B and its SoundMind-RL fine-tuned variant on the SoundMind benchmark under the text-to-audio setting.
+In this configuration, models receive text-only inputs and generate audio outputs.
+Results are reported using Word Error Rate (WER) to assess speech fidelity and accuracy (%) to measure reasoning correctness.
+SoundMind-RL significantly improves reasoning performance in the text-to-audio setting.
+As shown in Table 5, Qwen2.5-Omni-7B fine-tuned with SoundMind-RL achieves an accuracy of 83.84%, yielding a 3.05% improvement over the Qwen2.5-Omni-7B baseline.
+These results clearly indicate that the reward mechanism effectively encourages the model to produce coherent and consistent spoken reasoning, even without audio input.
+The observed accuracy gain is accompanied by a higher WER (6.99% vs.
+2.18%), which can be attributed to the generation of longer and more detailed reasoning sequences under reinforcement learning, thereby increasing the chance of recognition errors while preserving semantic correctness.
+5.3.3 Audio-to-Audio Reasoning
 […]
-In fully speech-based settings, SoundMind generalizes reasoning without textual anchors.
-Table 6 investigates the most demanding configuration: generating spoken reasoning from audio-only inputs, with no access to textual content during inference.
-In this dual-audio scenario, our model maintains a strong performance gain, raising accuracy from 77.59% to 81.40%.
-This demonstrates that SoundMind can internalize abstract logical dependencies in audio alone—without relying on textual shortcuts.
-Such robustness is especially valuable for real-world applications involving spoken dialogue systems, voice-based tutoring agents, or accessibility tools that require full audio-audio reasoning capabilities.
-Across all tasks, SoundMind consistently improves logical accuracy while maintaining acceptable speech quality.
-Our Qwen2.5-Omni-7B-RL model outperforms all baselines in reasoning accuracy across the three evaluated input-output modality pairs.
-These results confirm that reward-guided reinforcement learning can strengthen logical inference capabilities in audio-language models.
-While we observe a moderate rise in Word Error Rate (WER) in speech-based outputs, this trade-off is expected given the model’s increased focus on semantic correctness and structured explanation.
-The observed WER remains within acceptable bounds and does not hinder interpretability.
-Future work can explore fluency-aware RL objectives or dual-path reward balancing to further harmonize reasoning fidelity and audio naturalness.
-Model WER (%$\downarrow$ Acc. (%$\uparrow$
 Qwen2.5-Omni-7B 2.23 77.59
-Qwen2.5-Omni-7B-RL (Our) 8.95 81.40
-[…]
+Qwen2.5-Omni-7B (SoundMind-RL) 8.95 81.40
+Table 6: Performance of Qwen2.5-Omni-7B and its SoundMind-RL fine-tuned variant on the SoundMind benchmark under the audio-to-audio reasoning setting.
+In this configuration, models receive audio-only inputs and generate audio reasoning outputs.
+In the fully audio-based setting, our SoundMind-RL framework delivers notable performance gains under the most challenging condition.
+Table 6 presents results where models must derive and verbalize reasoning solely from acoustic input.
+Qwen2.5-Omni-7B fine-tuned with SoundMind-RL achieves 81.40% accuracy, improving substantially over the baseline (77.59%) and showing that the model can extract task-relevant logical dependencies directly from audio signals.
+The result is accompanied by a higher WER (8.95% vs.
+2.23%), which we attribute to the generation of longer, more elaborate reasoning sequences and the added difficulty of learning from raw acoustic input.
+Our analysis attributes this to the absence of explicit fluency constraints during reinforcement learning, which instead prioritizes structured reasoning and semantic correctness.
+Takeaway.
+Across all evaluated settings, SoundMind-RL consistently enhances logical reasoning performance for audio–language models.
+Qwen2.5-Omni-7B fine-tuned with SoundMind-RL achieves substantial gains across audio-to-text, text-to-audio, and audio-to-audio configurations, outperforming all baseline models.
+These results demonstrate that reward-guided reinforcement learning can reliably improve multimodal reasoning accuracy by encouraging faithful answer formatting, factual correctness, and sufficiently detailed reasoning traces.
+For speech-based outputs, we observe a moderate increase in WER relative to the baseline, which correlates with the production of longer, information-rich reasoning segments and suggests that the model prioritizes semantic completeness over surface fluency.
+Future work can explore fluency-aware or prosody-sensitive reward components to better balance reasoning fidelity with the naturalness of generated speech.
 6 Conclusion
-We introduce SoundMind, a novel rule-based reinforcement learning framework that empowers large-scale audio-language models with advanced logical reasoning capabilities across both audio and textual modalities.
-To enable such training, we build the Audio Logical Reasoning (ALR) dataset, a dual-modality benchmark comprising 6,446 high-quality samples annotated with chain-of-thought reasoning in both audio and text forms.
-Experimental results demonstrate that our method significantly improves performance and establishes new state-of-the-art results on the ALR benchmark, consistently outperforming strong baselines across three reasoning settings: text-to-audio, audio-to-text, and audio-to-audio.
-We hope this work can serve as a foundation for future research in logic-oriented large-scale audio-language modeling and in reinforcement-driven multimodal reasoning.
+We introduce SoundMind-RL, a novel rule-based reinforcement learning framework that empowers large-scale audio-language models with advanced logical reasoning capabilities across both audio and textual modalities.
+To support such training, we construct SoundMind, a dataset for the task of audio logical reasoning, containing 6,446 carefully curated audio–text pairs, each annotated with modality-specific chain-of-thought reasoning.
+Experimental results demonstrate that our method significantly improves performance and establishes state-of-the-art results on the SoundMind benchmark, consistently outperforming strong baselines across three reasoning settings: text-to-audio, audio-to-text, and audio-to-audio.
+We hope this work provides a useful resource and framework for advancing research in reasoning-oriented audio–language modeling and reinforcement-based multimodal learning.
 Limitations
-While our approach demonstrates strong performance, certain aspects warrant further exploration.
-In particular, the Word Error Rate (WER) in audio generation remains competitive, but still leaves room for improving reasoning fidelity during speech synthesis.
-Addressing this limitation could lead to more robust, adaptive, and human-aligned audio reasoning systems in future work.
+While this study shows promising results, several limitations remain.
+(1) The rule-based reward design enforces reasoning format consistency but may introduce rigidity, and its generalization to more diverse or open-ended tasks remains to be examined.
+(2) The dataset relies on synthetic speech and automatically generated chain-of-thought annotations; despite careful curation, subtle artifacts or biases may still be present and could influence model behavior.
+(3) The increase in Word Error Rate (WER) observed in audio generation suggests a possible trade-off between reasoning depth and fluency, motivating further work on prosody modeling and decoding strategies.
+We leave these challenges for future work, aiming to advance the development of audio reasoning systems that are more robust, generalizable, and aligned with human needs.
 Ethical Considerations
@@
 This study is supported by the Department of Defense grant HT9425-23-1-0267.
+Appendix A Baselines
+  • MiniCPM-o Yao et al.
+(2024): An open-source multimodal language model supporting vision, speech, and real-time streaming.
+It emphasizes efficiency and deployability, making it suitable for interactive multimodal understanding and generation across a variety of general-purpose and resource-constrained scenarios.
+  • Gemini-Pro-V1.5 Team et al.
+(2024): A multimodal model developed by Google that processes text, image, audio, and video inputs in a unified way.
+It is designed to handle extended context reasoning and cross-modal alignment, enabling coherent understanding and response in complex, multi-step tasks across modalities.
+  • Baichuan-Omni-1.5 Li et al.
+(2025): A multimodal model that combines text, vision, and audio understanding with audio generation.
+It aims to enable coherent cross-modal interaction and supports a wide range of multimodal and multitask applications in interactive settings.
+  • Qwen2-Audio Chu et al.
+(2024): A speech language model that accepts both audio and text as input for natural voice-based interaction.
+It can perform detailed audio analysis, transcription, and multilingual understanding, serving as a general-purpose audio–text reasoning system.
+  • Qwen2.5-Omni-7B Xu et al.
+(2025): A multimodal model that handles text, image, audio, and video inputs and generates both text and speech outputs.
+It supports streaming interactions and aims to provide robust multimodal reasoning and natural communication in real-time applications.
+Appendix B Text-to-Speech Model
+Constructing a fully human-recorded version of the audio logical reasoning dataset—covering thousands of user prompts, chain-of-thought reasoning traces, and final answers in both input and output modalities—would require extensive scripting, careful speaker coordination, and thousands of hours of professional narration.
+Such a large-scale recording effort is prohibitively expensive and logistically infeasible given the diversity and duration of reasoning sequences we target.
+Consistent with established practice in speech and multimodal learning research Jia et al.
+(2022); Zevallos (2022); Regmi et al.
+(2019); Cheng et al.
+(2024); Bartelds et al.
+(2023); Zhong et al.
+(2022), we synthesize the audio modality using a high-fidelity text-to-speech (TTS) system.
+This strategy enables scalable and reproducible generation of paired input–output audio segments, while preserving fine control over speaking rate, prosody, and acoustic quality—key factors for training models to perform multi-minute reasoning without performance degradation.
+To ensure that the synthesized audio faithfully reflects the complexity of logical reasoning, we conducted a systematic evaluation of several state-of-the-art open-source TTS systems.
+We compared MegaTTS-3 Jiang et al.
+(2025), WhisperSpeech OpenAI (2024), Spark-TTS Wang et al.
+(2025), and Zonos Zyphra (2024) across multiple dimensions, including perceived naturalness, prosody stability over long contexts, noise robustness, and intelligibility of extended reasoning chains.
+Each candidate system was benchmarked by synthesizing representative samples from our dataset and evaluated via multi-rater human listening studies.
+MegaTTS-3 consistently achieved the highest mean opinion scores, particularly in naturalness and prosodic consistency, qualities that we consider essential for faithfully preserving the chain-of-thought signal.
+Configuration $\lambda_{1}$ (TextFmt) $\lambda_{2}$ (AudioFmt) $\lambda_{3}$ (Answer) $\lambda_{4}$ (TextLen) $\lambda_{5}$ (AudioLen) Accuracy (%) $\Delta$ vs. full (pp)
+(1) w/o audio rewards 1.0 0 2.0 1.0 0 70.82 -10.58
+(2) w/o text rewards 0 1.0 2.0 0 1.0 48.84 -32.56
+(3) w/o answer reward 1.0 0.5 0 1.0 0.75 60.24 -21.16
+(4) full (Ours) 1.0 0.5 2.0 1.0 0.75 81.40
+Table 7: Ablation of the composite reward in SoundMind-RL, evaluated on the test set.
+$\lambda_{1}/\lambda_{2}$ enforce text/audio format compliance, $\lambda_{3}$ supervises answer correctness, and $\lambda_{4}/\lambda_{5}$ encourage reasoning length in text/audio space.
+Accuracy is reported in %, and $\Delta$ indicates percentage-point ($p$) difference relative to the full model.
+On the basis of these findings, we select MegaTTS-3 as the synthesis engine for both user prompts and model reasoning traces.
+Its ability to produce long, coherent, and acoustically stable speech makes it well suited for the extended reasoning sequences characteristic of the ALR task.
+The resulting corpus provides a controlled yet realistic training signal that aligns with our objective of enabling robust, audio-native logical reasoning in large audio-language models.
+We further acknowledge the open-source speech community for providing high-quality TTS systems, which have been instrumental in enabling reproducible, large-scale construction of the audio datasets.
+Appendix C Potential Applications
+Reasoning-capable ALMs can enable applications that demand explainability, transparency, and natural interaction.
+We highlight key domains where audio–native reasoning offers clear value and outline open challenges for future research.
+Education and Tutoring.
+Spoken step-by-step explanations can help learners understand complex problems by exposing intermediate reasoning rather than presenting only a final answer Wang et al.
+(2020).
+This is particularly relevant for subjects such as logic, mathematics, and music, where reasoning over multimodal signals is essential for mastery.
+Recent studies on music-related question answering and multimodal reasoning further highlight the need for models that can integrate audio and symbolic information to support learning and feedback You et al.
+(2025); Diao et al.
+(2025).
+Accessibility and Screen-Free Interaction.
+For users with visual impairments, hands-busy scenarios, or clinical settings, verbal reasoning offers an inclusive interface to access complex information.
+Replayable and summarizable reasoning can support comprehension and patient-centered decisions, consistent with efforts on trustworthy medical response generation Li et al.
+(2024) and healthcare voice assistants Zhan et al.
+(2024).
+Ensuring natural prosody and coherence over multi-minute speech remains crucial for high-stakes applications.
+Conversational Agents and Voice Assistants.
+In applications such as troubleshooting, procedural guidance, and eligibility checking, users often ask “why” in addition to “what.” Models capable of structured, spoken justifications could reduce repeated follow-ups and improve user trust.
+Meetings, Lectures, and Knowledge Capture.
+Beyond transcription, audio–native reasoning models could verbalize intermediate inferences, summarize discussions, and surface entailment relations in real time, turning speech streams into structured knowledge.
+Robustness to overlapping speech, noise, and speaker variability remains a key challenge, as exemplified by the AMI Meeting Corpus Carletta et al.
+(2005).
+Appendix D Ablation Analysis
+Quantitative impact.
+Table 7 reports the results of ablating individual reward components.
+Removing text-format/length supervision (Row 2) causes the most severe degradation, with accuracy dropping to 48.84% (-32.56% relative to the full model).
+This sharp decline underscores that structural guidance on the text side is essential for producing well-formed and logically verifiable reasoning traces.
+Excluding the answer-correctness reward (Row 3) reduces accuracy to 60.24% (-21.16%), indicating that explicit factual supervision is necessary to ensure generated reasoning converges to the correct entailment decision rather than drifting into irrelevant but superficially coherent explanations.
+Removing audio-format/length rewards (Row 1) still lowers accuracy to 70.82% (-10.58%), suggesting that modality-consistent formatting and duration constraints regularize the shared policy and enhance robustness.
+Overall, these results demonstrate that each reward term contributes meaningfully to model performance, and that their combination is crucial for maximizing reasoning accuracy.
+Key takeaways.
+Table 7 highlights three main takeaways: (1) Text-side format and length supervision is the dominant factor for producing reliable, verifiable reasoning traces.
+(2) Answer-correctness reward is essential to align reasoning with ground truth labels and convert fluent explanations into correct decisions.
+(3) Audio-side rewards serve as powerful regularizers, improving policy robustness and benefiting reasoning quality across modalities.